What the letters s h t o l h a spell?
Well to begin with, thing of words that have those letters in them. Such as:
1. hat
2. holt
3. hot
4. as
5. ash
It could be a name of someone:
Sholtha
Athlosh
Tolhash
Now look at the words and decide. Could it be this? Or could it be this? Maybe that? Give yourself a challenge and mix them up a bit. If it doesn't make a word, it will most likely make a name! Have fun! :)
Is tossing a coin and spinning a spinner dependent?
Tossing a coin and spinning a spinner are independent events. The outcome of one does not affect the outcome of the other; the coin can land on heads or tails regardless of the spinner's result, which can land on any of its designated sections. Therefore, the probability of each event remains unchanged regardless of the outcome of the other.
How do you get identical twins and fraternal twins?
Identical twins are two children born of one egg and one sperm, that splits early on and then separates, each half then developing into a full organism. Since they came from only one egg and one sperm, their genes will be identical.
Fraternal twins are two children born of two eggs and two sperms, each of them developing into a full organism. Since they did not come from only one egg and one sperm, their genes will be different. They could even be opposite sexes.
What is the distribution with a mean of 0 and a standard deviation of 1?
It is any standardised distribution.
Note:
In listing the contents of the marble jar, you mentioned yellow marbles twice ...
"2 yellow marbles" and "1 yellow marble". We'll assume there are 3 yellow marbles.
Also note:
You didn't actually say that the numbers 1 through 10 are printed on the 10 sides
of the 10-sided die. We'll assume that they are.
Also note:
You didn't actually say that we select a marble at the same time that we toss the
10-sided die. We'll assume that we're supposed to do that.
Now that we've done the proof-reading and cleaned up the question, we can go ahead
and answer it:
The probability of rolling a 10 with the 10-sided die is 1 out of 10 = 0.1 .
There are 18 marbles total in the jar, and 3 of them are yellow. The probability of
drawing one yellow marble out of the jar is (3/18) = 1/6 .
The probability of drawing 1 yellow marble and rolling a 10 on the die is
(1/6) x (0.1) = 12/3%
What is the probability of tossing a coin 5 times and getting all tails?
The probability of tossing a coin 5 times and getting all tails is:
P(TTTTT) = (1/2)5 = 0.03125 ≈ 3.13%
What is the chance of drawing a heart in a standard deck of cards?
The probability is 12/52 or 1/4 since 12 of the 52 cards are hearts.
Explain in detail the three possible outcomes that may result from this research?
The question is very poorly specified so this answer is simply a wild guess at what the questioner might want.
Three possible outcomes of any research, designed to test some hypothesis, are:
(a) evidence in support of the hypothesis;
(b) evidence disproving the hypothesis; or
(c) evidence that can neither prove (support) nor disprove the hypothesis.
When you boil something what happens?
A chemical change occurs when you boil something. Also, evaporation can occur when it gets hot enough.
How is using survey one way to determine experimental probability?
is bye it must be an even # of people but ramdom people so its fair cause other people do like differnt thing so pike pepole who are different
Distinguish between selective distribution and exclusive distribution?
Selective distribution involves a producer using a limited number of outlets in a geographical area to sell products. An advantage of this approach is that the producer can choose the most appropriate or best-performing outlets and focus effort (e.g. training) on them. Selective distribution works best when consumers are prepared to "shop around" - in other words - they have a preference for a particular brand or price and will search out the outlets that supply.
Exclusive distribution is an extreme form of selective distribution in which only one wholesaler, retailer or distributor is used in a specific geographical area.
What happens in the event of a vacancy occuring in either chamber of congress?
they call for a special election to fill the vacancy.
What is an example of a result that would have a probability near 0 percent?
All life on earth dying out by the end of this year.
There is.
According to Benford's law the probability that the leading digit of a measure is x is approximately
log10(1 + 1/x)
The law appears in nature: lengths of rivers, heights of mountain peaks and so on. This suggests that small events are much more likely than larger ones.
For more on Benford's law, see the link.
Assuming the die is a standard die with a different number from the set {1, 2, 3, 4, 5, 6} on each side, then:
Probability_of_success = number_of_ways_of_success/total_number_of_outcomes
There is only 1 way to roll a 4 and there are 6 possible outcomes, therefore the probability of rolling a 4 is 1/6
The die has no knowledge of previous rolls; each roll is independent, thus:
The probability of three fours is a row is the probability of a 4 times the probability of a 4 times the probability of a 4, which is:
Probability = 1/6 × 1/6 × 1/6 = 1/216
Go with options...a.56 b.86 c.24 d.120 now i choose 24...ratio is 3:4:5..divide 24 by 3, 4 and 5,,, by these third man will get 10 rs..so the ans is 24
How many possible outcomes are there when flipping a coin 5 times?
There are 25 = 32 possible outcomes.